OpenStudy (anonymous):

if f(x)= 4-x and g(x)=x-4, then g(f(x^2))=

OpenStudy (anonymous):

=-x^2

OpenStudy (anonymous):

f(x^2)= 4-x^2 g(f(x^2) = 4-x^2 - 4 = -x^2